Halley's comet, which passes around the Sun every 76 years, has an elliptical orbit. When closest to the Sun (perihelion) it is at a distance of 8.823×10^10m and moves with a speed of 54.6km/s. When farthest from the Sun (aphelion) it is at a distance of 6.152×10^12mand moves with a speed of  783m/s.

Find the angular momentum of Halley's comet at perihelion. (Take the mass of Halley's comet to be 9.8×10^14kg.)

Express your answer using two significant figures.

Lp

Lp

=

nothing

  kg⋅m2/s  

SubmitRequest Answer

Part B

Find the angular momentum of Halley's comet at aphelion.

Express your answer using two significant figures.

La

La

=

nothing

  kg⋅m2/s  

Please help ASAP I'm not sure what to do thank you :)
Answer #1

Part A.

Angular momentum is given by:

L = m*v*r

for Comet at perihelion,

m = mass of Halley's comet = 9.8*10^14 kg

v = speed at perihelion = 54.6 km/sec = 54600 m/sec

r = radius at perihelion = 8.823*10^10 m

Using these values:

L = 9.8*10^14*54600*8.823*10^10

L = 4.72*10^30 kg-m^2/sec

Part B.

for Comet at aphelion,

m = mass of Halley's comet = 9.8*10^14 kg

v = speed at aphelion = 783 m/sec

r = radius at aphelion = 6.152*10^12 m

Using these values:

L = 9.8*10^14*783*6.152*10^12

L = 4.72*10^30 kg-m^2/sec

Please Upvote.